Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night report 2021-04-09

Observers: Priscila Pessi
Support: Christian Vogl, Andreas Floers, Geza Csoernyei, Alexander Holas and Stefan Taubenberger
TAT: Giuliano Pignata, Jacob Teffs, Michel Dennefeld

EFOSC Calibrations: 
Bias, dome_flats, arcs.
Skyflats  UBVR (last R flats are not usefull because the stars appeared)

Weather:
Start: humidity 51%, wind   4.9 m/s, seeing 1.3
Middle: humidity 18%  wind  7.3  m/s, seeing 1.05
End: humidity 21% wind 6.4  m/s, seeing 1.22

Comments:
23:13 UT image analysis.
05:00 UT, CCD had a problem the OB was aborted although the integration was finished.

EFOSC

 UTTarget 
 TypeSetup SlitExposure (s)Airmass
Seeing 
Notes
23:38
AT2021blz
followup
gr13
1.5
2700
1.4
1.65

00:31
SN2020abtf
followup
gr13
1
2700
1.5
1.34
variable seeing during integration.
 01:21l745a
std
gr11,13,16
1,1.5
-
1.2
1.53

01:41
SDSS_J0938_1253
followup
gr11,gr16
1
1500 each
1.4
1.24

 02:45 SN2021dbg followup gr11 1 1500 1.2 1.24 
 03:18 AT2021iff classificationgr13
 1 4501.5
 1.15DIMM has stopped, value of seeing comes from the quality of the acquisition image. Type Ia, z= 0.024
 03:33SN2021zj
followup
gr11
 12700
 2.01.36
Value of seeing comes from the quality of the acquisition image.
 04:25 SN2021gno followup gr11 1 1800
1.4
 0.9DIMM active again. The CCD showed and error and the OB was aborted when the integration finished. TIO told me they will generate a ticket to see if the data can be recovered.
 05:08 SN2021gno followup gr13 1 1800 1.5 1.DIMM has stopped, value of seeing comes from the quality of the acquisition image.
 05:40 gd153 std gr11,13,16 1,1.5 - 1.7 1.1 DIMM active again.

SOFI: Switch at 05:58 UT, image analysis and focus after switch.
One of the quadrants was a bit noisy, we did our best to reduce the noise to the maximum possible.

UT
Target
Type
Setup
Exposure
Airmass
Seeing
Notes
06:24
SN2021fxy
followup
JH
00:05:03
1.1
1.02

06:37
gd153
std
BG
00:05:42
2.0
0.98

06:43
Hip080118
telluric
BG
00:04:43
1.9
0.94

 06:51 AT2017gge followup BG 01:01:10 1.7 0.91 
 08:16 AT2017gge followup JHK 00:33:39 1.7 0.91 
08:46
SN2021hwk
followup
JH
00:52:44
1.5
0.91
There was a lot of noise in one quadrant so we aborted, move the rotator and started again.
 09:44 SN2021hem followup JH 00:17:29 1.6 1.2 
Comments